I need help answering f(x)=x^2+3 and g(x)=-x+9 find(g-f)(-8)
Answer by Alan3354(69443)

I need help answering f(x)=x^2+3 and g(x)=-x+9 find(g-f)(-8)
---------------
That's f(-8) - g(-8)
---
f(x) = x^2+3
Sub -8 for any x
f(-8) = (-8)^2 + 3 = 67
---
g(x) = -x + 9
g(-8) = -(-8) + 9 = 17
----
(g-f)(-8) = 17 - 67 = -50